OpenStudy (anonymous):

I have a math word problem that is killing me. A pole is painted red, blue and green. 1/3 is painted red. 1/2 is painted blue. 4 feet is painted green. What is the height of the pole.

OpenStudy (anonymous):

So, we write this out as: \[ 4=\left(1-\frac{1}{3}-\frac{1}{2}\right)P=\frac{P}{6} \]So, therefore, \(\frac{1}{6}\) of the pole is green, so, by this: \[ 4\cdot6=P=24 \]So the pole is 24 ft.

OpenStudy (anonymous):

1/3 is red, 1/2 is blue, remainder (three feet) is green. The part that isn't painted green = red + blue = 1/3 + 1/2 = 2/6 + 3/6 = 5/6 So part that is painted green = 1 - 5/6 = 1/6 And 1/6 = 3 feet, so 6/6 = 18 feet 1 = 18 feet. So length of pole is 18 feet.

OpenStudy (anonymous):

Instead of 3, it would be 4, but good explanation.

OpenStudy (anonymous):

Thank both of you this was a lesson learned...LolWolf thanks for showing me the formula